IMACS progress report no. 55
B. Bigelow 6 November 2002
I) Project Status
Currently scheduled tasks: Status:
a) DSS drive assembly/testing (Hare, Kowal) IP
b) DSS grating tilt mech. testing (Hare) Done
c) DSS new grating tilt mech. des. (Hare) Done
d) DSS new grating tilt drafting (Cleary) Done
e) DSS new grating tilt fab. (Kowal) IP
f) DSS new grating cell design (Bigelow) Done
g) DSS new grating cell drafting (Cleary) Done
h) DSS new grating cell fab (Kowal) IP
i) Calibration system mech. des. (Hare) on hold
j) CCD array assembly (Thompson/Burley) IP
k) Grism procurement (Sutin) IP
l) Grism cell drafting (Cleary) Done
m) Grism cell fabrication (Kowal) IP
n) Entrance hatch fabrication (Kowal) IP
o) Short Camera shutter assembly (Hare) ready to start
p) Instrument electronics assembly (Bagish) on sched.
q) Control software design (Birk) on sched.
r) Short Camera opto-assembly (Bigelow) IP
s) Instrument enclosures installation (Gunnels) Done
t) Instrument utility wrap fabrication (Rettig) Done
u) Instrument utility wrap installation Done
v) IFU design and build (U. Durham) IP
Project events in November:
- Short camera aspheres finished at Tinsley and shipped to STF for AR coating
- Grating flexure testing completed, new cell and tilt mechanisms in progress
- Collimator, field lens, slit mask, and long camera alignment completed to spec.
- Enclosure integration now complete
- Electronics cables and utility wrap mostly completed
Latest project images:
Final Assy: http://www.ociw.edu/instrumentation/imacs/devel/final/index.html
II) Optics (Sutin)
* Tinsley finished and delivered our aspheres to Spectrum Thin Films.
A report is available on Tinsley's performance, by request.
* Spectrum TF has, as of Tuesday, coated S01 and was about to coat
* S04-mark-II is still due on Nov 15.
* The Dewar window returned from Spectrum TF with a water-resistant
* 300 l/mm grisms is at Harold Johnson being re-polished. Due to be
finished this week.
* 600 l/mm grisms is at Harold Johnson being glued. Due to be
finished this week.
* The echellette prism size has been computed. I need to hand over
the final size to Bruce to see if it fits.
III) Electronics (Bagish)
- Instrument end manifold attached.
- All MCS panels engraved and all except one panel installed.
- All except for five wrap cables are complete.
- Wire up individual subsections.
- Attach RTD thermal sensors.
- PI Enclosure: select enclosure, mount components, wire and plumb.
IV) Software (Birk/Clardy)
- Gratings/Disperser Wheel: We successfully tested the grating control
loop. For November I plan to continue testing of the grating
mechanism and start running the disperser wheel.
- SH-, PR- and CF-Guiders: The guiders were re-assembled and all
functions (motors, limits, etc) successfully tested. For November I
plan to run the guiders extensively while rotating the structure.
- DSP/Science Array Software: Using the MIKE chips as a test bed for
IMACS we learned a lot about how to read and transfer images from the
DSP over the CCD-server to the user-level software. For early
November we plan to read all 8 chip simultaneously for the first
- 2nd shutter: no progress (not assembled)
- Mask Server: We tested all functions (motors, limits, sensors) and
all but one position sensor work. We tested the repeatability of the
mask-transfer mechanism it it performed well. During November I
would like to stress-test the mask server, but that requires the
removal of an interference between the magnetic latch and the mask
- A simulator program was constructed to accept guider movement
commands and reply with guider position and status. Using this, the
"GMAP" program was extended to control the IMACS guiders and show
their locations and movement on the display. The movement of
guiders is initiated by the "P1" or "P2" buttons in GMAP, consistent
with their use in moving other guide probes. The projection
geometry of GMAP was changed to a vector method, giving improved
accuracy and a consistent inverse location of mouse click location
to celestial position.
- Some more design was done for the Astronomer's GUI (interface to
mask generation input data).
- Mask cutting was improved to allow a complex shape to an alignment
hole; this is for a test mask in which each hole is surrounded by 4
slits. The "hole with 4 slits" is defined as a complex shape and is
cut at each location using the set of hole dimensions. Also, output
formats were cleaned up to make the .nc file more readable.
- Discussions with Pat McCarthy regarding the "nod & shuffle" mode of
observation indicate that adding one parameter to the slit
dimensions will suffice to support this mode. Such a change looks
to be straightforward, but has not yet been implemented.
V) Detector Systems (Thomson/Burley)
1) Re-assembled the IMACS array, attached the two refurbished
cryotigers, pumped and cooled the array, and read out 7 of 8 chips.
CCD "B" is the troublesome chip, and had problems previously when
read through the other output amplifier. We need to check the "B"
preamp and clock signals to the chip.
The operating temperature of the array is about 10 degC warmer than
expected. The chip temperatures range from -85 to -92 degC. The
dark current is about 4 e-/pixel/hour at this temperature.
2) We did extensive testing of the red+blue MIKE ccds at LCO, managed
to clear up some quirks in the system, and test in mock-operation
with Christoph's control software.
2a) Revised the optical link code in the Altera chips on the PCI
interface and the dsp boards. Changed the scrambler code to
17-bits (from 15-bits). The optical link uses a self-synchronous
scrambler, which has the potential to occasionally freeze. This
results in bursts of errors at the receive end of the link.
Christoph adjusted his code to be somewhat error tolerant. The
17-bit scrambler is only a partial solution -- some errors still
result, but are now fairly rare. A 21-bit or 23-bit scrambler is a
better solution, but will have to be shoe-horned into the Altera
2b) got rid of some low-level bias instability (horizontal banding) by
adding some extra NOPs during the preamp settling periods.
Changing the dual-slope trim circuit cleared up some injected noise
and bias tilt which was low enough not to notice in high gain mode,
but showed up noticeably in low-gain mode. Low-gain bias
instability has been a persistent problem, so getting rid of it was
2c) tested the MIKE dewars in operation on the nasmyth platform, with
the telescope drives operating and the dome rotating. There is no
apparent interference with the ccd readout.
2d) revised the dsp PID code and tested it. The PID code is able to
control the ccd temperature to a few tenths of a degree C.
2e) changed the base.asm code to use NOPs for timing the dual-slope
integration periods (rather than timers). Data is now sent by dma
from the dsp (in preparation for 8-chip readout).
3) a sample of the latest read noise measurements (from the MIKE ccds,
2100x4096 pixels, with 1x1 binning and high gain/low gain
Read Time Pixel time e/DN rms e/DN rms
--------- ---------- ---- --- ---- ---
63 sec 6.2 usec 2.02 6.5 e- 4.91 9.2
85 sec 8.8 usec 1.06 4.7 e- 2.56 5.7
130 sec 13.6 usec 0.54 3.6 e- 1.31 4.0
170 sec 19.0 usec 0.37 3.2 e- 0.88 3.4
VI) Mechanics (Bigelow/Hare)
1) Continued short camera opto-mechanics assembly
2) Alignment and testing of slit mask stability completed
3) Grism cell fab drawings to shop, fab in progress
4) Ordered final cryotiger hose assemblies for both cameras
5) Completed new grating cell design, fab in progress
6) Fabrication of hatch assembly in progress
7) Grism cell fabrication in progress
8) Designed FSS insertion stages mounts for both cameras,
fab in progress
9) Ordered, received, and assembled IGUS chain for utility wrap
10) Made some progress on instrument cooling system plumbing
1) Finish short camera assembly - ready for testing by end of Nov.
2) Assemble grism cells if fab permits
3) Assemble instr. hatch of fab permits
4) Assemble grating cell in new tilt mechanism for testing
5) Install filter servers for both cameras with new mounts
6) Progress cooling system design
* Continued opto-mechanical and electro-mechanical assembly and testing
of all Guide Cameras.
* Stripped, anodized, and reassembled all guide cameras.
* Designing counterbalance for the Shack-Hartmann Guide Camera.
* Redesigned Grating Tilt Mechanism trunion assembly. Ordered
associated parts. Fabrication in progress.
* Reviewing design of counterbalance for GTM.
* Assembly and testing of DSS Drive and Cable Wrap.
* Modifying Clamp Assembly and latch mechanisms.
* Continue assembly, testing, and rework of all guide cameras as necessary.
* Continued testing of the Grating Tilt Mechanism.
* Assembly and testing of the DSS Drive Assembly and Cable Wrap Subassembly.
* Assembly of the second shutter.
* Resume design work of the Calibration Unit.
VII) Fabrication (Kowal)
1) Finished Carousel Drive Assembly
2) Hatch is 80% finished (drilling on cover/base plates remain)
3) Brackets, Trusses & Truss Base for Grism Cells are finished.
4) All guiders have been anodized.
5) New aluminum grating cell and base are finished.
6) Stainless encoder sector for G.T.M. is finished.
7) Stainless drive sector for G.T.M. is 90% finished.
8) Stainless trunion for G.T.M. is currently in progress.
Expectations for November:
1) Finish all new G.T.M. Parts.
2) Finish Hatch.
3) Have Bulkhead Shroud fabricated.
4) Work on Grism Cells.
No report this month.
IX) Other Business
Next meeting: 10:15 AM, Thursday (!), 12/5/02